# Role of flow topology in wind-driven wildfire propagation

### v1

@@ 4.2.1 Role of wind oscillation - St Number
To effectively investigate the interaction between transient wildfires and time-varying wind topology, we select a wind oscillation timescale $(1/\Omega)$ that aligns with the total fuel reaction timescale $\overline{t}_{r}$. A transient wildfire simulation with a steady double-gyre wind velocity $(\Omega=0)$ determined the fuel consumption timescale $\overline{t}_{r}\sim 10^{-2}$. Consequently, we select the wind oscillation frequency to be on the order of $\overline{t}_{r}$. In Fig. 6, the spatio-temporal evolution of the Heaviside firefront at various time instants is presented under an unsteady double-gyre wind flow, oscillating at a frequency of $\textit{St}=2$, with an oscillation amplitude of $\lambda=0.25$ and a velocity magnitude of $A_{m}=0.1$. Figure 6a illustrates the fuel and temperature fields along with the corresponding instantaneous rLCS and aLCS results. At $\tau=0.001$, the initialized Heaviside firefront is advected left, right, and downward, resulting in three distinct firefronts–the left firefront $F_{L1}^{X}$, the right firefront $F_{R1}^{X}$, and the bottom firefront $F_{B}^{Y}$–with dotted lines indicating their locations within the temperature field. Upon examining the fuel bed traces, it becomes evident that the left firefront advances considerably further than the right firefront. This is attributed to the influence of the aLCS-1, which specifically drives the advection of the left firefront while exerting no effect on the right firefront. In contrast, the right firefront experiences moderated advection due to the presence of rLCS-2, which repels and decelerates its progression. Although aLCS-2 is anticipated to contribute to the advancement of the right firefront, its influence is diminished by the preceding rLCS-2, which constrains the progress of the right firefront. This deceleration effect is most pronounced while rLCS-2 remains in the upstream region of the right firefront. However, since wildfire transport is not a pure advection problem and LCS cannot form transport barriers, ultimately, the firefront passes rLCS-2, and aLCS-2 begins to accelerate the advection of the right firefront, as observed at $\tau=0.0025$. The deceleration action exerted by rLCS-2 is distinctly evident in the tracked instantaneous location plot of $F_{R1}^{X}$, illustrated in Fig. 6b, where the red box highlights a notable reduction in the instantaneous firefront’s group velocity, followed by a subsequent increment corresponding to the traversal of rLCS-2. This differential behavior emphasizes the contrasting roles of aLCS-2 and rLCS-2 in shaping the firefront dynamics. Moreover, the bottom firefront does not advance uniformly downward but is distorted due to the decelerating influence of rLCS-1, as evidenced by the non-parallel alignment of the identified dotted line in the temperature field.
@@
Figure 7: Spatio-temporal evolution of the Heaviside firefront under a double gyre wind velocity field, characterized by parameters $A_{m}=0.1$, $\textit{St}=7.5$, and $\lambda=0.25$, over a uniformly distributed fuel bed, computed at $\Phi=10^{-3}$, $\textit{Da}=10^{3}$, and $\epsilon=0.03$. (a) The fuel and temperature fields at various time instants, along with the corresponding repelling and attracting LCS plots, with streamlines superimposed over the fuel bed. (b) The instantaneous spatial advection of the four firefronts is plotted. (c) The FTLE field corresponding to the time-averaged wind velocity fields superimposed with time-averaged streamlines.
@@
Next, the transient behavior of wildfires is investigated under an unsteady double gyre flow oscillating at a high frequency, characterized by a Strouhal number $\textit{St}=7.5$, employing the same other parameters, including an oscillation amplitude of $\lambda=0.25$ and a velocity magnitude of $A_{m}=0.1$. This case aims to elucidate the wildfire dynamics when the time scales of fuel consumption and wind flow oscillation are disparate. Figure 7 illustrates the spatio-temporal evolution of the Heaviside firefront under the considered unsteady double gyre flow scenario As anticipated, the advection of all firefronts nearly exhibits symmetric advancement across the fuel bed, since the flow oscillation time is considerably smaller than the fuel consumption time scale, implying that the firefronts do not have sufficient time to respond to the wind oscillations. Consequently, the subsequent emergence of top firefronts (resulting from the advection of right/left firefronts) also displays a nearly symmetric advection. The instantaneous locations of the tracked firefronts, as depicted in Fig. 7a, further substantiate this nearly symmetric advection ($t_{left}\approx t_{right}$) over the evolution time $\tau$. Although changes in local group velocity for both the right and left firefronts are evident, as highlighted within the red box panel, the local deceleration/acceleration is minimal compared to the last transient case. The instantaneous FTLE fields and corresponding aLCS and rLCS reveal an asymmetric pattern, which appears to be unrelated to the high-Strouhal number oscillation scenario. Instead, FTLE calculation on the time-averaged wind velocity field, as shown in Fig. 7c is sufficient to characterize the advection of firefronts in this case. Notably, the LCS manifests as a ridge-like structure along the dividing streamline. In the present case, all firefronts adhere to the time-averaged local streamlines, much like the behavior observed in a saddle-type steady wind velocity scenario.
By comparing two Strouhal number cases, it becomes evident that asymmetric advection of firefronts is more pronounced when $\textit{St}=2$, where wind oscillation closely aligns with the fuel consumption time scale. In contrast, minimal asymmetry is observed when $\textit{St}=7.5$, where the respective time scales are misaligned, indicating that the degree of asymmetry is most notable when $\textit{St}\rightarrow 1$. To address this quantitatively, one can project the advection characteristics of the firefronts onto a Bode plot–similar to the approach used by Ducruix et al. (2000) for transient premixed flames–thereby providing a comprehensive understanding of the wildfire’s response to unsteady wind conditions.
@@ 4.2.2 Dynamic response of wildfires to unsteady wind oscillation
The dynamic response of wildfires to the imposed unsteady double-gyre wind topology is systematically quantified by projecting their advection characteristics onto a Bode plot through a transfer function approach. In the Bode plot, a transfer function (TF) is generally defined as $TF=\overline{R}\angle\phi$, where magnitude $(\overline{R})$ denotes the magnitude gain, and the phase angle $(\phi)$ indicates the phase difference of an output and an input response.
@@
It is important to note that the magnitude of the TF quantifies the degree of asymmetry between the advection of the right and left firefronts by calculating their time-averaged group velocity, thereby reflecting the uneven spread of the fire under the influence of the unsteady double-gyre wind. A TF magnitude different from unity indicates asymmetric transport, suggesting that similar asymmetry should be observed in the advection of subsequent formations of other top firefronts, as discussed earlier. Additionally, the phase angle of the TF shows the phase response of the firefronts relative to a baseline scenario characterized by a steady wind topology ($\textit{St}=0$), where a negative phase angle indicates a phase lag and a positive phase angle denotes a phase advance in the response of the firefront progression to wind oscillation. Figure 8 presents the Bode plot illustrating the transient wildfire characteristics of firefronts subjected to an unsteady double gyre flow, demonstrating the transfer function of both right and left firefronts across varying Strouhal numbers and oscillation amplitudes, $\lambda\in(0.05,0.15,0.25)$. The $\overline{R}_{L}$ plot is omitted as it is the inverse of $\overline{R}_{R}$. Notably, the phase response of both firefronts is not monotonous, reaching zero for certain Strouhal numbers, which indicates the existence of a phase inversion frequency in the transient wildfire propagation scenario under unsteady wind conditions, implying a phase-switching phenomenon where the fire response shifts from advance to lag or vice versa, as similarly observed in Motta et al. (2015); Viknesh and Poddar (2021) for unsteady flow control applications.
Figure 8: Bode plots – $(a)$ The magnitude of the $TF_{R}$, $(b)$ the phase angle of the TF for the right firefront, and $(c)$ the phase angle of the TF for the left firefront are shown. These results are computed with $A_{m}=0.1$, $\Phi=10^{-3}$, $Da=10^{3}$, and $\epsilon=0.03$ for different wind oscillation amplitudes $(\lambda)$ and Strouhal numbers.
@@ 5 Conclusion
In the present research, the wildfire transport model proposed by Asensio and Ferragut (2002) has been employed to study the role of flow topology under two benchmark wind flow typologies: steady wind with saddle-type fixed points and unsteady wind characterized by double gyre flow. The key outcomes of this study are enumerated as follows:
@@
$(c)$ A robust wildfire transport solver has been developed, leveraging CUDA support within a finite difference method framework. The solver employs compact schemes and implicit-explicit Runge-Kutta methods to resolve both spatial and temporal solutions while incorporating the LAD model to selectively attenuate high-frequency instabilities without compromising accuracy in smooth regions.
@@
Future research should focus on several key areas to advance our understanding of wildfire dynamics. Incorporating realistic vegetation distributions, including firebreaks, along with two-way coupled wind velocity computations using the Navier-Stokes equations and wildfire equations, will provide a more nuanced perspective on wind-fire interactions. Additionally, integrating actual terrain topography, including both uphill and downhill scenarios, is crucial for accurately capturing the effects of terrain on wildfire behavior. Examining the relevance of LCS within these more complex and realistic settings will enhance predictive models and deepen our comprehension of complex wildfire propagation patterns.
@@ Acknowledgment
The developed wildfire transport python solver will be available on GitHub after peer review.
Appendix A Implicit-Explicit time integration algorithm
IMEX-RK scheme for the new wildfire model
Initial solution $\text{Initial solution }\boldsymbol{T}^{n}\text{ and }\boldsymbol{\beta}^{n}\text{ at }\tau=\tau^{n}$
do $\text{do }i=1,\ldots,s$

### v2

@@ 4.2.1 Role of wind oscillation - St Number
To effectively investigate the interaction between transient wildfires and time-varying wind topology, we select a wind oscillation timescale $(1/\Omega)$ that aligns with the total fuel reaction timescale $\overline{t}_{r}$. A transient wildfire simulation with a steady double-gyre wind velocity $(\Omega=0)$ determined the fuel consumption timescale $\overline{t}_{r}\sim 10^{-2}$. Consequently, we select the wind oscillation frequency to be on the order of $\overline{t}_{r}$. In Fig. 6, the spatio-temporal evolution of the Heaviside firefront at various time instants is presented under an unsteady double-gyre wind flow, oscillating at a frequency of $\textit{St}=2$, with an oscillation amplitude of $\lambda=0.25$ and a velocity magnitude of $A_{m}=0.1$. Fig. 6a illustrates the fuel and temperature fields along with the corresponding instantaneous rLCS and aLCS results. At $\tau=0.001$, the initialized Heaviside firefront is advected left, right, and downward, resulting in three distinct firefronts–the left firefront $F_{L1}^{X}$, the right firefront $F_{R1}^{X}$, and the bottom firefront $F_{B}^{Y}$–with dotted lines indicating their locations within the temperature field. Upon examining the fuel bed traces, it becomes evident that the left firefront advances considerably further than the right firefront. This is attributed to the influence of the aLCS-1, which specifically drives the advection of the left firefront while exerting no effect on the right firefront. In contrast, the right firefront experiences moderated advection due to the presence of rLCS-2, which repels and decelerates its progression. Although aLCS-2 is anticipated to contribute to the advancement of the right firefront, its influence is diminished by the preceding rLCS-2, which constrains the progress of the right firefront. This deceleration effect is most pronounced while rLCS-2 remains in the upstream region of the right firefront. However, since wildfire transport is not a pure advection problem and LCS cannot form transport barriers, ultimately, the firefront passes rLCS-2, and aLCS-2 begins to accelerate the advection of the right firefront, as observed at $\tau=0.0025$. The deceleration action exerted by rLCS-2 is distinctly evident in the tracked instantaneous location plot of $F_{R1}^{X}$, illustrated in Fig. 6b, where the red box highlights a notable reduction in the instantaneous firefront’s group velocity, followed by a subsequent increment corresponding to the traversal of rLCS-2. This differential behavior emphasizes the contrasting roles of aLCS-2 and rLCS-2 in shaping the firefront dynamics. Moreover, the bottom firefront does not advance uniformly downward but is distorted due to the decelerating influence of rLCS-1, as evidenced by the non-parallel alignment of the identified dotted line in the temperature field.
@@
Figure 7: Spatio-temporal evolution of the Heaviside firefront under a double gyre wind velocity field, characterized by parameters $A_{m}=0.1$, $\textit{St}=7.5$, and $\lambda=0.25$, over a uniformly distributed fuel bed, computed at $\Phi=10^{-3}$, $\textit{Da}=10^{3}$, and $\epsilon=0.03$. (a) The fuel and temperature fields at various time instants, along with the corresponding repelling and attracting LCS plots, with streamlines superimposed over the fuel bed. (b) The instantaneous spatial advection of the four firefronts is plotted. (c) The FTLE field corresponding to the time-averaged wind velocity fields is superimposed with time-averaged streamlines.
@@
Next, the transient behavior of wildfires is investigated under an unsteady double gyre flow oscillating at a high frequency, characterized by a Strouhal number $\textit{St}=7.5$, employing the same other parameters, including an oscillation amplitude of $\lambda=0.25$ and a velocity magnitude of $A_{m}=0.1$. This case aims to elucidate the wildfire dynamics when the time scales of fuel consumption and wind flow oscillation are disparate. Fig. 7 illustrates the spatio-temporal evolution of the Heaviside firefront under the considered unsteady double gyre flow scenario As anticipated, the advection of all firefronts nearly exhibits symmetric advancement across the fuel bed, since the flow oscillation time is considerably smaller than the fuel consumption time scale, implying that the firefronts do not have sufficient time to respond to the wind oscillations. Consequently, the subsequent emergence of top firefronts (resulting from the advection of right/left firefronts) also displays a nearly symmetric advection. The instantaneous locations of the tracked firefronts, as depicted in Fig. 7a, further substantiate this nearly symmetric advection ($t_{left}\approx t_{right}$) over the evolution time $\tau$. Although changes in local group velocity for both the right and left firefronts are evident, as highlighted within the red box panel, the local deceleration/acceleration is minimal compared to the last transient case. The instantaneous FTLE fields and corresponding aLCS and rLCS reveal an asymmetric pattern, which appears to be unrelated to the high-Strouhal number oscillation scenario. Instead, FTLE calculation on the time-averaged wind velocity field, as shown in Fig. 7c is sufficient to characterize the advection of firefronts in this case. Notably, the LCS manifests as a ridge-like structure along the dividing streamline. In the present case, all firefronts adhere to the time-averaged local streamlines, much like the behavior observed in a saddle-type steady wind velocity scenario.
By comparing two Strouhal number cases, it becomes evident that asymmetric advection of firefronts is more pronounced when $\textit{St}=2$, where wind oscillation closely aligns with the fuel consumption time scale. In contrast, minimal asymmetry is observed when $\textit{St}=7.5$, where the respective time scales are misaligned, indicating that the degree of asymmetry is most notable when $\textit{St}\rightarrow 1$. To address this quantitatively, one can project the advection characteristics of the firefronts onto a Bode plot–similar to the approach used by Ducruix et al. (2000) for transient premixed flames–thereby providing an understanding of the wildfire’s response to unsteady wind conditions.
@@ 4.2.2 Dynamic response of wildfires to unsteady wind oscillation
The dynamic response of wildfires to the imposed unsteady double-gyre wind topology is systematically quantified by projecting the firefronts advection characteristics onto a Bode plot through a transfer function approach. In the Bode plot, a transfer function (TF) is generally defined as $TF=\overline{R}\angle\phi$, where magnitude $(\overline{R})$ denotes the amplitude gain, and the phase angle $(\phi)$ indicates the phase difference of an output and an input response.
@@
It is important to note that the magnitude of the TF quantifies the degree of asymmetry between the advection of the right and left firefronts by calculating their time-averaged group velocity, thereby reflecting the uneven spread of the fire under the influence of the unsteady double-gyre wind. A TF magnitude different from unity indicates asymmetric transport, suggesting that similar asymmetry should be observed in the advection of subsequent formations of other top firefronts, as discussed earlier. Additionally, the phase angle of the TF shows the phase response of the firefronts relative to a baseline scenario characterized by a steady wind topology ($\textit{St}=0$), where a negative phase angle indicates a phase lag and a positive phase angle denotes a phase advance in the response of the firefront progression to wind oscillation. Fig. 8 presents the Bode plot illustrating the transient wildfire characteristics of firefronts subjected to an unsteady double gyre flow, demonstrating the transfer function of both right and left firefronts across varying Strouhal numbers and oscillation amplitudes, $\lambda\in(0.05,0.15,0.25)$. The $\overline{R}_{L}$ plot is omitted as it is the inverse of $\overline{R}_{R}$. Notably, the phase response of both firefronts is not monotonous, reaching zero for certain Strouhal numbers, which indicates the existence of a phase inversion frequency in the transient wildfire propagation scenario under unsteady wind conditions, implying a phase-switching phenomenon where the fire response shifts from advance to lag or vice versa, as similarly observed in Motta et al. (2015); Viknesh and Poddar (2021) for unsteady flow control applications.
Figure 8: Bode plots – $(a)$ The magnitude of the $TF_{R}$, $(b)$ the phase angle of the $TF_{R}$, and $(c)$ the phase angle of the $TF_{L}$ are shown. These results are computed with $A_{m}=0.1$, $\Phi=10^{-3}$, $Da=10^{3}$, and $\epsilon=0.03$ for different wind oscillation amplitudes $(\lambda)$ and Strouhal numbers.
@@ 5 Conclusion
In the present research, the reactive flow-wildfire transport model proposed by Asensio and Ferragut (2002) has been employed to study the role of flow topology under two benchmark wind flow typologies: steady wind with saddle-type fixed points and unsteady wind characterized by double gyre flow. The key outcomes of this study are enumerated as follows:
@@
$(c)$ A robust wildfire transport solver has been developed, leveraging CUDA support within a finite difference method framework. The solver employs the upwind compact schemes and implicit-explicit Runge-Kutta methods to resolve both spatial and temporal solutions while incorporating the LAD model to selectively attenuate high-frequency instabilities without compromising accuracy in smooth regions.
@@
Although extensive numerical wildfire simulations were conducted, it is important to recognize that the findings of this study are derived from a simplified wildfire model that focuses on reactive flow dynamics, which was specifically chosen to investigate the fundamental relevance of LCS to wind-driven wildfire propagation by resolving only the flame scale. Additionally, the wildfire model assumes that the radiative flux only acts locally by treating the medium as optically thick, adopting the Rosseland approximation, wherein the radiative heat flux is expressed as a nonlinear diffusion. Alternatively, this assumption can be relaxed by employing a global radiation model, as proposed in Margerit and Séro-Guillaume (2002); Ferragut et al. (2007), which accounts for long-range radiative interactions and, in a more recent method, Navas-Montilla et al. (2024) incorporates coupled global radiative–wind effects, albeit with increased computational complexity. Nevertheless, the correspondence of stable and unstable manifolds of wind topology with firefront progression is analyzed within this simplified, constrained framework, it may still serve as a viable basis for extending its applicability to real-world wildfire scenarios. Hence, future research should adopt a more comprehensive approach by employing fully coupled CFD-wildfire methodologies to further advance the understanding of wildfire dynamics and the correspondence of LCS in more complex and realistic settings. In addition, incorporating detailed vegetation distributions, including firebreaks, and accounting for actual terrain topography—such as both uphill and downhill scenarios—is essential for accurately capturing terrain effects on wildfire behavior. Ultimately, these additional effects will not only enhance predictive models but also provide deeper insights into the intricate mechanisms governing wind-driven wildfire propagation.
@@ Acknowledgment
The developed GPU-accelerated Wildfire Transport FDM Python Solver is publicly available on GitHub: https://github.com/siva-viknesh/Wildland-Fire-Dynamics.
Appendix A FDM Solver Algorithm
This section presents the FDM solver algorithm developed to simulate wildfire propagation, incorporating the selected physics-based CDR combustion model, Asensio and Ferragut (2002). The numerical solver employs the optimized fifth-order upwind compact scheme (OUCS2) for the convection term and a second-order central difference scheme (CD2) for both physical diffusion and localized artificial diffusion components. Temporal integration is performed using the IMEX-RK method, which enables simultaneous integration of stiff ($\overline{\boldsymbol{T}}^{(i)}$) and non-stiff ($\widehat{\boldsymbol{T}}^{(i)}$) temperature and fuel fields. Two wind topology scenarios are considered: (i) a steady wind field characterized by a saddle-type fixed-point structure, and (ii) a transient wind field governed by a double-gyre configuration, introducing time-varying wind dynamics parameterized by the Strouhal number. The complete FDM solver algorithm is detailed in Algorithm 1.
For the steady wind case, the algorithm begins by initializing the temperature and fuel fields, denoted as $\boldsymbol{T}^{0}(x)$ and $\boldsymbol{\beta}^{0}(x)$, respectively. Additionally, physical and numerical parameters, relevant non-dimensional numbers, and the wind velocity field are prescribed. At each time step $n$, the solver first enforces the boundary condition representing zero-flux constraints at the domain boundaries. Subsequently, for each IMEX-RK stage $i$, the temperature field is partitioned into an explicit (non-stiff) component $\widehat{\boldsymbol{T}}^{(i)}$ and an implicit (stiff) component $\overline{\boldsymbol{T}}^{(i)}$, while the fuel field $\boldsymbol{\beta}^{(i)}$ is updated implicitly. At each temporal stage $i$, the slope $F_{i}^{\beta}$ associated with the fuel field is computed via a closed-form expression obtained from the discretized stiff fuel equation. The temperature slope $F_{i}^{T}$ is then evaluated by summing the following four contributions:
  • Convection ($\mathcal{C}_{i}^{n}$): Evaluated using the OUCS2 operator, this term captures the advective transport of firefronts driven by the prescribed wind field and is treated explicitly.
  • Diffusion ($\mathcal{D}_{i}^{n}$): Incorporates both stiff and non-stiff temperature components. The thermal diffusivity $K(\widehat{\boldsymbol{T}}^{(i)})$ is computed from the non-stiff temperature component, while spatial derivatives of the stiff temperature are approximated using the CD2 scheme. Here, the first derivative form of the CD2 scheme is employed twice, one after the other, due to its conservative form representation.
  • Reaction and Natural Convection ($\mathcal{R}_{i}^{n}$): Represents the nonlinear source and sink terms. The reaction term arising from fuel consumption is treated implicitly, while the natural convection component is handled explicitly.
  • Localized Artificial Diffusion ($\mathcal{L}_{i}^{n}$): The spatially varying artificial diffusion coefficient $\mu^{n}$ is determined using a biharmonic operator acting on the non-stiff temperature, followed by a Heaviside function that isolates regions near the firefront. This coefficient is scaled by a tunable parameter $C_{\mu}$. Then the localized artificial term is computed by applying a Laplacian operator to the non-stiff temperature, scaled by $\mu^{n}$.
The total temperature slope $F_{i}^{T}$ is then constructed by accounting for all four contributions and is solved using the iterative Newton–Raphson method with a convergence tolerance of $10^{-12}$. Here, $a_{ij}$ and $\tilde{a}_{ij}$ represent the coefficients in a standard Butcher tableau, used to calculate internal temporal stages of the IMEX-RK scheme setting. Upon completing all temporal stages, the temperature and fuel fields are updated through a weighted combination of the computed stage slopes, thereby advancing the solution in time. For the transient wind topology case, the only modification to the procedure involves the computation and update of the instantaneous wind velocity field characterized by the Strouhal number, at each temporal marching.
Algorithm 1 Wildfire CDR Algorithm
Input:
 Initial fields: $\boldsymbol{T}^{0}(x)$, $\boldsymbol{\beta}^{0}(x)$
 Time step: $\Delta\tau$, number of steps: $N_{\tau}$
 Parameters: $\overline{\kappa}$, $\alpha$, $\epsilon$, $q$, $\overline{T}_{pc}$, $C_{\mu}$
 Non-dimensional numbers: $\Phi$, $Da$
 Temporal stages: s
 Velocity field: $\overrightarrow{\mathbf{w}}(\xi,\eta)$
Operator Construction:
 First derivative operators: $\nabla$ (OUCS2), $\overline{\nabla}$ (CD2)
 Laplacian: $\nabla^{2}$ (CD2), Biharmonic: $\nabla^{4}$ (CD2)
Temporal Marching:
for $n=0$ to $N_{\tau}-1$ do
     Apply Boundary Condition : $\left[\overrightarrow{\mathbf{w}}\overline{T}^{n}-K(\overline{T}^{n})\nabla\overline{T}^{n}\right]\cdot\hat{n}=0$
     Compute stage values for $2\leq i\leq s$:
  $\widehat{\boldsymbol{T}}^{(i)}\leftarrow\boldsymbol{T}^{n}+\Delta\tau\displaystyle\sum_{j=1}^{i-1}\tilde{a}_{ij}F_{j}^{T}$,  $\overline{\boldsymbol{T}}^{(i)}\leftarrow\boldsymbol{T}^{n}+\Delta\tau\displaystyle\sum_{j=1}^{i-1}a_{ij}F_{j}^{T}$,  $\boldsymbol{\beta}^{(i)}\leftarrow\boldsymbol{\beta}^{n}+\Delta\tau\displaystyle\sum_{j=1}^{i-1}a_{ij}F_{j}^{\beta}$
     Solve for $F_{i}^{\beta}$:
$F_{i}^{\beta}\leftarrow-\frac{\epsilon}{q}\frac{\boldsymbol{\beta}^{(i)}s(\widehat{\boldsymbol{T}}^{(i)})^{+}\exp(\widehat{\boldsymbol{T}}^{(i)}/[1+\epsilon\widehat{\boldsymbol{T}}^{(i)}])}{\left[1+(\epsilon/q)\Delta\tau a_{ii}s(\widehat{\boldsymbol{T}}^{(i)})^{+}\exp(\widehat{\boldsymbol{T}}^{(i)}/[1+\epsilon\widehat{\boldsymbol{T}}^{(i)}])\right]}$
     Solve a linear system for each $F_{i}^{T}$:
    Convection:
$\mathcal{C}_{i}^{n}\leftarrow\frac{\overrightarrow{\mathbf{w}}}{\Phi}\cdot\nabla\widehat{\boldsymbol{T}}^{(i)}$
    Diffusion:
$\mathcal{D}_{i}^{n}\leftarrow\frac{1}{Da}\,\overline{\nabla}\cdot\left[K(\widehat{\boldsymbol{T}}^{(i)})\,\overline{\nabla}\left(\overline{\boldsymbol{T}}^{(i)}+\Delta\tau a_{ii}F_{i}^{T}\right)\right]$
    Reaction and Natural Convection:
$\qquad\mathcal{R}_{i}^{n}\leftarrow s(\widehat{\boldsymbol{T}}^{(i)})^{+}\left[\boldsymbol{\beta}^{(i)}+\Delta\tau a_{ii}F_{i}^{\beta}\right]\exp(\widehat{\boldsymbol{T}}^{(i)}/[1+\epsilon\widehat{\boldsymbol{T}}^{(i)}])-\alpha\left[\overline{\boldsymbol{T}}^{(i)}+\Delta\tau a_{ii}F_{i}^{T}\right]$
    Localized Artificial Diffusion:
$\mathcal{L}_{i}^{n}\leftarrow\mu^{n}\nabla^{2}\widehat{\boldsymbol{T}}^{(i)};\quad\mu^{n}\leftarrow C_{\mu}\,\mathbf{H_{F}}\left[|\nabla^{4}\widehat{\boldsymbol{T}}^{(i)}|\right]$
    Total slope at stage $i$:
$\qquad F_{i}^{T}\leftarrow\mathcal{D}_{i}^{n}+\mathcal{R}_{i}^{n}-\mathcal{C}_{i}^{n}+\mathcal{L}_{i}^{n}$
end for
  $\boldsymbol{T}^{n+1}\leftarrow\boldsymbol{T}^{n}+\Delta\tau\displaystyle\sum_{j=1}^{s}b_{j}F_{j}^{T}$,  $\boldsymbol{\beta}^{n+1}\leftarrow\boldsymbol{\beta}^{n}+\Delta\tau\displaystyle\sum_{j=1}^{s}b_{j}F_{j}^{\beta}$
 Marched solution: $\boldsymbol{T}^{n+1}$, $\boldsymbol{\beta}^{n+1}$ at $\tau^{n+1}=\tau^{n}+\Delta\tau$
Appendix B Numerical FDM solver validation
In this section, we validate our solver using a benchmark 2D perfect initial Heaviside firefront over a uniform fuel distribution scenario, as presented in Bürger et al. (2020). Resolving the advecting Heaviside firefront over a long integration time poses a significant numerical challenge, as it excites all possible frequencies, including spurious ones associated with Gibbs’ oscillation in a Fourier spectrum. Despite this challenge, the OUCS2 compact scheme adeptly captures the spatiotemporal evolution of the Heaviside firefront effectively, even without incorporating the LAD model for the considered wildfire scenario (Fig. 9). We maintain consistency by adhering to the parameters in Bürger et al. (2020): a square domain size of $50$, $\epsilon=3\times 10^{-2}$, $q=1$, $\alpha=10^{-3}$, $\overline{T}_{\text{pc}}=3$, $\overline{\kappa}=0.1$, initial firefront temperature $\overline{T}=31$, wind speed $|\overrightarrow{\mathbf{w}}|=425$ with a direction of $45^{\circ}$, $C_{\mu}=0$, and $Da=\Phi=1$. Furthermore, with the initial firefront precisely positioned at coordinates $(8.0,8.0)$ and a square side length of 4, and the fuel uniformly initialized at 0.6 across the spatial domain, the simulation is conducted with a time step of $\Delta\tau=10^{-7}$ over a grid size of $256\times 256$. The results depicted in Fig. 9, illustrate the spatio-temporal evolution of the Heaviside firefront at different time instants, demonstrating good agreement with the Bürger et al. (2020) results. Discrepancies may arise from using a more accurate IMEX-RK3 time integration and OUCS2 schemes in our work, contrasting with the employment of IMEX-RK2 scheme and highly dissipative WENO scheme in Bürger et al. (2020).
Appendix C Methodological Comparison of Wildfire Models
In this section, a comparative analysis is presented between three wildfire modeling approaches: the semi-empirical model introduced by Rothermel (1972), the CDR wildfire model utilized in this work, and the fully coupled wildfire CFD method, FIRETEC. The comparison is conducted by examining the rate of spread ($R_{s},m/s$) predictions for a benchmark wildfire scenario involving fireline propagation over a flat terrain under unidirectional wind conditions. Specifically, the performance of the CDR wildfire model is evaluated against the study of Pimont et al. (2012), which provides a comparison between FIRETEC simulations and the Rothermel model across various slope and wind conditions. Mathematically, the semi-empirical Rothermel model expresses the rate of spread as
$R_{s}=R_{0}(1+\phi_{u}+\phi_{s})\;,$ (20)
where $R_{0}$ denotes the rate of spread under no-wind and zero-slope conditions, $\phi_{u}$ accounts for the influence of wind (upwind or downwind), and $\phi_{s}$ represents the contribution due to terrain slope. In Pimont et al. (2012), the simulations were conducted within a cuboidal computational domain with a square base of $320\times 320$ m, a vertical extent of 615 m, comprising 60 simulations, spanning three wind velocities (1, 5, 12 m/s), two initial fireline lengths (20 m and 50 m), and 10 different slope conditions. Their results indicated that the dependence of spread rate on different wind and slope conditions is neither purely multiplicative nor strictly additive, deviating from the semi-empirical formulation given by Eq. 20. Furthermore, a nonlinear coupling between wind, slope, and fireline length was identified as a key factor influencing the rate of spread. For the Rothermel model calculation, the coefficient value of $R_{0}$ and the wind-related coefficient $\phi_{u}$ were computed using the FIRETEC simulation data. In contrast, the slope-related coefficient $\phi_{s}$ was evaluated using the expression given in Rothermel (1972).
For the present comparison, we consider Quercus coccifera, one of the fuels analyzed in Pimont et al. (2012), whose properties are summarized in Table 2. The rate of spread ($R_{s}$) results are extracted from Fig. 2a of Pimont et al. (2012), which presents the comparative predictions across varying wind and slope conditions; however, to remain consistent with the primary focus of this comparative analysis, only the zero-slope cases ($\phi_{s}=0$) and three wind velocity scenarios for initial fireline length of 50 m are selected. Our CDR wildfire simulations are configured to closely align with the three-dimensional FIRETEC setup, with model parameters and non-dimensional groups computed using the mean values of physical quantities, as listed in Table 2. It should be noted that such averaging introduces an independent modeling problem, which may affect predictive accuracy.
     Material
     $\rho$
     (kg/m3)
     $C$
     (kJ/kg·K)
     $k$
     (W/m·K)
     $A$
     (s-1)
     $\epsilon$
     Quercus coccifera      700      1.3–1.5      0.12–0.20      $10^{7}$$10^{12}$      0.016–0.038
     Air      1.293      1.005      0.025–0.030      –      –
     Mean values      100      1      0.1      $10^{10}$      0.03
Table 2: Material properties employed in the FIRETEC wildfire simulations, as reported in Pimont et al. (2012) and the references therein, along with the representative mean values adopted in the present comparative analysis.